OpenStudy (anonymous):

how to calculate: 1+1/1+1/1x2+1/1x2x3

OpenStudy (nincompoop):

try to write it properly by using parenthesis and brackets and square brackets

OpenStudy (anonymous):

can we use eulers formula?

OpenStudy (anonymous):

1+(1/1)+(1/1x2)+(1/1x2x3)

OpenStudy (nincompoop):

it's up to you. you can even use thermodynamics or hawking's radiation if you want, but what good are they to you if you're not even writing your expressions appropriately?

OpenStudy (anonymous):

i just did

OpenStudy (nincompoop):

it looks like just a normal expression with fractions … why were you thinking of Euler's Formula?

OpenStudy (anonymous):

because there is another question related to the above using eulers formula

OpenStudy (nincompoop):

so? Simplify the following: 1+1/1+1/2+1/(2×3) One raised to any number is one. 1/1 = 1: 1+1+1/2+1/(2×3) Multiply 2 and 3 together. 2×3 = 6: 1+1+1/2+1/6 Put the fractions in 1+1+1/2+1/6 over a common denominator. Put 1+1+1/2+1/6 over the common denominator 6. 1+1+1/2+1/6 = 6/6+6/6+3/6+1/6: 6/6+6/6+3/6+1/6 Add the fractions over a common denominator to a single fraction. 6/6+6/6+3/6+1/6 = (6+6+3+1)/6: (6+6+3+1)/6 Evaluate 6+6+3+1. 6+6+3+1 = 16: 16/6 Reduce 16/6 to lowest terms. The gcd of 16 and 6 is 2, so 16/6 = (2×8)/(2×3) = 2/2×8/3 = 8/3

OpenStudy (anonymous):

what if the calculation continued forever

OpenStudy (anonymous):

1+(1/1)+(1/1x2)+(1/1x2x3)+(1/1x2x3x4)+... +...+...

OpenStudy (anonymous):

@satellite73